# AlgoPuzzleVQA

### 06 Mar 2024 v1

@@ 2.1 Setup and Baselines
We perform all our experiments on a multi-choice question-answering setup. We create three negative answer choices for each instance by using heuristics such as randomly sampling numbers within the same magnitude as the gold answer. More details can be found in LABEL:sec:appendix:mcqa. In total, we thus have four answer choices - one gold positive and three random negatives, for all puzzles except one. The exception is the Board Tiling puzzle, where we have Yes and No as the possible choices. We evaluate various closed and open-source multimodal language models on our dataset. We consider the following models: GPT-4V gpt4vision and Gemini Pro geminipro as the closed models; InstructBLIP Vicuna 7B and 13B Dai2023InstructBLIPTG, and LLaVA-1.5 13B liu2023improved as the open-sourced models. We use the accuracy of predicting the final answer as the evaluation metric.
@@ 2.3 Main Results
GPT-4V Gemini Pro GPT-4V Gemini Pro
@@
  Colour 35.0 32.8 34.4 31.4 31.2 29.2 32.1
  Position 30.3 30.2 31.7 28.1 29.1 27.6 29.1
  Shape/Size 23.9 25.6 27.6 27.1 29.8 27.8 27.2
  Text 33.2 28.5 31.8 28.0 25.7 28.0 27.3
@@
  Arithmetic 30.3 30.2 31.7 28.1 29.1 27.6 29.1
  Boolean Logic 27.6 29.4 32.1 29.8 31.1 29.1 31.4
  Combinatorics 37.5 33.0 33.0 34.5 22.0 30.0 30.5
  Graphs 33.6 32.2 35.0 29.4 27.2 29.6 28.4
  Optimization 25.6 28.1 27.0 25.2 29.1 28.1 25.2
  Search 26.3 28.4 28.1 26.4 27.8 26.1 25.7
  Sets 31.0 33.0 34.3 30.7 30.0 31.0 34.7

### 10 Mar 2024 v2

@@ 2.1 Setup and Baselines
We perform all our experiments on a multi-choice question-answering setup. We create three negative answer choices for each instance by using heuristics such as randomly sampling numbers within the same magnitude as the gold answer. More details can be found in LABEL:sec:appendix:mcqa. In total, we thus have four answer choices - one gold positive and three random negatives, for all puzzles except one. The exception is the Board Tiling puzzle, where we have Yes and No as the possible choices. We evaluate various closed and open-source multimodal language models on our dataset. We consider the following models: GPT-4V gpt4vision, Gemini Pro geminipro, Claude 3 Opus [^5] https://www.anthropic.com/news/claude-3-family as the closed models; InstructBLIP Vicuna 7B and 13B Dai2023InstructBLIPTG, and LLaVA-1.5 13B liu2023improved as the open-sourced models. We use the accuracy of predicting the final answer as the evaluation metric.
@@ 2.3 Main Results
GPT-4V Gemini Pro Claude 3 GPT-4V Gemini Pro Claude 3
@@
 Colour 35.0 32.8 33.2 34.4 31.4 32.8 31.2 29.2 32.1
 Position 30.3 30.2 30.9 31.7 28.1 31.2 29.1 27.6 28.5
 Shape/Size 23.9 25.6 30.0 27.6 27.1 29.5 29.8 27.8 25.9
 Text 33.2 28.5 34.7 31.8 28.0 33.2 25.7 28.0 25.0
@@
 Arithmetic 30.3 30.2 30.9 31.7 28.1 31.2 29.1 27.6 28.5
 Boolean Logic 27.6 29.4 35.4 32.1 29.8 32.4 31.1 29.1 31.8
 Combinatorics 37.5 33.0 38.0 33.0 34.5 37.5 22.0 30.0 30.5
 Graphs 33.6 32.2 30.8 35.0 29.4 28.6 27.2 29.6 28.4
 Optimization 25.6 28.1 26.6 27.0 25.2 29.0 29.1 28.1 25.6
 Search 26.3 28.4 28.3 28.1 26.4 27.6 27.8 26.1 26.0
 Sets 31.0 33.0 39.7 34.3 30.7 40.7 30.0 31.0 34.7